a bathroom measures 250cm by 175cm.calculate the side of the largest square tile that can be used without cutting




Answer by Theo(13342)

the greatest common denominator is 25.

25 * 10 = 250
25 * 7 = 175

your biggest tile would be 25 * 25 = 625 square units.

you would need 10 * 7 = 70 of them.

total square units is 250 * 175 = 43750

70 * 625 = 43750.

that should do it.